# Encourage or Inhibit Monosemanticity? Revisit Monosemanticity from a Feature Decorrelation Perspective

### v1

@@ 5.1.1 Setup
[…]
@@
[…]
@@
We compare with DPO and SimDPO (Meng et al., 2024), which uses the average log probability of a sequence as the implicit reward and introduce a target reward margin to the to encourage a larger reward, i.e., $-\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)$.
[…]
@@ 5.1.2 Preference optimization results
In some cases, DPO is inferior to SFT, i.e., the Sycophancy dataset for Llama2-base. Similar issues are observed on SimDPO, it is inferior on both the CogReframe and Sycophancy datasets (the two smaller datasets) for Llama2-chat. This can be explained by the relatively limited data leading to model overfitting, a phenomenon theoretically and empirically detected in DPO (Azar et al., 2024). Instead, DecPO improves upon DPO performance due to its efficiency in decreasing the overfitting issue and is generally superior to SFT.
@@ 5.1.4 Effects of different layers
We study the effects of implementing the feature decorrelation regularizer in different layers, noting that the regularizer is applied to only one model block. The results for Llama2-7b-hf and Llama-2-7b-chat-hf can be seen in Figure 7. We observe that performance is highly sensitive to layer selection, which can be attributed to varying degrees of monosemanticity across layers. Interestingly, optimal results are not consistently observed in the last layers; instead, the middle layers are optimal for the Toxicity dataset, while for Cognition Reframing, the optimal layers are at very early stages. This suggests cumulative effects where constraints applied in earlier layers impact representations in deeper layers, as also observed in prior knowledge editing studies Meng et al. (2023).
@@ 5.2 Theoretical insights
$\vspace{-5mm}-\text{log}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right).$ (6)
@@
where $r(x,y)$ is the pointwise reward given by the LLMs, and $\sigma$ is a normalization term for the probability. Consider a simple example where $y$ is always preferred over $y^{\prime}$, i.e., $p(y^{\prime}-y|x)=1$. In this case, the model is driven to create a very high reward discrepancy $(r(y)-r(y^{\prime}))\rightarrow+\infty$, especially if there are limited preference data. In other words, ranking-based DPO tends to overfit on training samples to attain lower loss, which often leads to over-exploitation of shortcut features Geirhos et al. (2020) to hack the reward function (implicitly defined in DPO). Therefore, the proposed decorrelation regularization is an effective strategy to prevent such reward overfitting by encouraging the models to learn diverse features from the data. As shown previously, this regularizer also helps the model to learn more monosemantic features during training and enhance model interpretability.
@@ 6 Conclusion
In this paper, we have revisited recent studies in monosemanticity probing and proposed a monosemanticity proxy via feature decorrelation perspective. To study the research question Should monosemanticity be encouraged or inhabited in a model level for alignment training? we experimentally provide the empirical evidence that the alignment, such as DPO, can improve monosemanticity. We have also clarified that there is no clear relation between the monosemanticity degree and model size. Then, we have studied the effects of enhanced monosemanticity via applying a decorrelation regularizer in DPO training. We observed that representation diversity and activation sparsity are co-occurred, valid the effectiveness of our proxy. The evidence from the better alignment experiment further verifies our hypothesis that monosemanticity should be encouraged for better model capacity.
@@
However, much of our analysis is ad hoc, tailored to the specific feature being investigated, and requires substantial researcher effort to draw conclusions. While we explored models of varying sizes, they were all from the same llama family and trained with limited data. Additionally, the largest model we studied is llama3-8b, which is still more than an order-of-magnitude off the frontier. Given the emergent abilities of LLMs with scale, it is possible our analysis misses a key dynamic underlying the success of the largest models. Moreover, we do not give strict theory proof about the relation between monosemanticity and feature decorrelation, and we only use cosine similarity between every two different samples as an indication, without constructing input-feature dataset applied in monosemanticity probing.
@@ Acknowledgements
This work was supported in part by the UK Engineering and Physical Sciences Research Council (EPSRC) through a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2) and a New Horizons grant (grant no. EP/X019063/1).

### v2

@@ 5.1.1 Setup
We include three datasets covering different aspects of human values that existing LLMs should align with, i.e., Toxicity (Lee et al., 2024), Cognition Reframing (CogFrame (Sharma et al., 2023) and Sycophancy (Perez et al., 2022)  [^5] The dataset details are in Appendix A.1.
@@
We follow the practice of using advanced LLMs as evaluators, which demonstrates a high correlation with human evaluation (Wang et al., 2023). GPT-3.5 is provided with the criteria and generated outputs and is required to make a binary decision about whether the outputs align with the criteria  [^6] The prompt details are in Appendix A.2.
@@
We compare with DPO and SimDPO (Meng et al., 2024), which uses the average log probability of a sequence as the implicit reward and introduce a target reward margin to encourage a larger reward, i.e., $-\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)$.
Additionally, we compare with zero-shot in-context learning (ICL) and supervised fine-tuning (SFT). We include $\mathcal{L}_{1}$ regularization, which is commonly used to encourage activation sparsity. [^7] We also used ReLU as a sparsity enhancement by replacing the original SiLU activation in MLP with ReLU, but the model collapsed.
@@ 5.1.2 Preference optimization results
In some cases, DPO is inferior to SFT, i.e., the Sycophancy dataset for Llama2-base. Similar issues are observed on SimDPO, it is inferior on both the CogReframe and Sycophancy datasets (the two smaller datasets) for Llama2-chat. This can be explained by the relatively limited data leading to model overfitting, a phenomenon theoretically and empirically observed for DPO (Azar et al., 2024). Instead, DecPO improves upon DPO performance due to its efficiency in decreasing the overfitting issue and is generally superior to SFT.
@@ 5.1.4 Effects of different layers
We study the effects of implementing the feature decorrelation regularizer in different layers, noting that the regularizer is applied to only one model block. The results for Llama2-7b-hf and Llama-2-7b-chat-hf can be seen in Figure 8. We observe that performance is highly sensitive to layer selection, which can be attributed to varying degrees of monosemanticity across layers. Interestingly, optimal results are not consistently observed in the last layers; instead, the middle layers are optimal for the Toxicity dataset, while for Cognition Reframing, the optimal layers are at very early stages. This suggests cumulative effects where constraints applied in earlier layers impact representations in deeper layers, as also observed in prior knowledge editing studies (Meng et al., 2023).
@@ 5.2 Theoretical insights
$-\text{log}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right).$ (5)
@@
where $r(x,y)$ is the pointwise reward given by the LLMs, and $\sigma$ is a normalization term for the probability. Consider a simple example where $y$ is always preferred over $y^{\prime}$, i.e., $p(y^{\prime}-y|x)=1$. In this case, the model is driven to create a very high reward discrepancy $(r(y)-r(y^{\prime}))\rightarrow+\infty$, especially if there are limited preference data (Azar et al., 2024). In other words, ranking-based DPO tends to overfit on training samples to attain lower loss, which often leads to over-exploitation of shortcut features Geirhos et al. (2020) to hack the reward function (implicitly defined in DPO). Therefore, the proposed decorrelation regularization is an effective strategy to prevent such reward overfitting by encouraging the models to learn diverse features from the data. As shown previously, this regularizer also helps the model to learn more monosemantic features during training and enhance model interpretability.
@@ 6 Conclusion
In this paper, we have revisited recent studies in monosemanticity probing and proposed a monosemanticity proxy via feature decorrelation perspective. To study the research question Should monosemanticity be encouraged or inhabited in a model level for alignment training? we experimentally provide the empirical evidence that the alignment, such as DPO, can improve monosemanticity. We have also clarified that there is no clear relation between the monosemanticity degree and model size. Then, we have studied the effects of enhanced monosemanticity via applying a decorrelation regularizer in DPO training. The evidence from the better alignment experiment further verifies our hypothesis that monosemanticity should be encouraged for better model capacity.
@@
However, much of our analysis is ad hoc, tailored to the specific feature being investigated, and requires substantial researcher effort to draw conclusions. While we explored models of varying sizes, they were all from the same llama family and trained with limited data. Additionally, the largest model we studied is llama3-8b, which is still more than an order-of-magnitude off the frontier. Given the emergent abilities of LLMs with scale, it is possible our analysis misses a key dynamic underlying the success of the largest models.
@@ Acknowledgements
This work was supported in part by the UK Engineering and Physical Sciences Research Council (EPSRC) through a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2) and a New Horizons grant (grant no. EP/X019063/1), and Innovate UK through the Accelerating Trustworthy AI programme (grant no. 10093055). Y. Wang is supported by Office of Naval Research grant N00014-20-1-2023 (MURI ML-SCOPE).